in this series 6,4,1,2,2,8,7,4,2,1,5,3,8,6,2,2,1,4,1,3,5,8,6,how many pairs of successive numbers have a difference of 2 each ?

Answer by Edwin McCravy(20056)

Underline them:

6,4,1,2,2,8,7,4,2,1,5,3,8,6,2,2,1,4,1,3,5,8,6

Answer: There are five. (6,4), (4,2), (5,3), (8,6) and (8,6)  

Note: We did not underline (1,3) or (3,5) because their 
difference is 1-3 = 3-5 = -2, not 2.

To find the difference between two numbers, it's always 
"first minus second", never "second minus first".

Edwin